OpenStudy (anonymous):

Solve -4(x + 10) - 6 = -3(x - 2).

OpenStudy (anonymous):

need help please

OpenStudy (anonymous):

\(\Large -4(x + 10) - 6 = -3(x - 2).\) \(\Large (-4)x+(-4)10-6=(-3)x+(-3)(-2)\) \(\Large-4x-40-6=-3x+6\) \(\Large 4x-3x=-40-6-6\) \(\Large x=-52\)

OpenStudy (anonymous):

so 52 wrong right that what i got at first.

OpenStudy (anonymous):

yes \(\Large x=52\) is wrong The correct is \(\Large x=-52\)

OpenStudy (anonymous):

ok thank you and do you think you could help me if just one more please Solve for x. -ax + 2b > 8

OpenStudy (anonymous):

\(\Large -ax+2b>8\) \(\Large -ax>8-2b\) \(\Large ax<2b-8\) \(\Large x<\frac{2b-8}{a}\)

OpenStudy (anonymous):

wow thank you so much you help me a lot

OpenStudy (anonymous):

\(\Large \color{blue}{Ur}\ \color{green}{Welcome}\ \color{red}{=)} \)